OpenStudy (anonymous):

Would (1/8)^-2 simplify to 8^2? or 8/1^2?

OpenStudy (lalaly):

8^2

OpenStudy (anonymous):

1 / (1/8)^2 = 8^2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

With negative exponents, change the base to its reciprocal, and then make the exponent positive. Example: 3^(-1/2) = (1/3) ^(1/2)